OpenStudy (anonymous):

The slope of the line whose equation is 3y + 2x = 1 is as an answer i got -2/3?

OpenStudy (anonymous):

Mlhuyck, I posted a solution for that problem we were working on

OpenStudy (anonymous):

i got the answer already :)

OpenStudy (anonymous):

Your answer here is good. 3y = 1 - 2x y = 1/3 - 2/3x y = mx + b, with m representing slope, so -2/3 is correct.